OpenStudy (anonymous):

Need help on how to simplify (8s^2-4s+12)/(s(s^2+4)) to (5 s-4)/(s^2+4) + 3/s

OpenStudy (anonymous):

laplace?

OpenStudy (anonymous):

yeah, I'm trying to do that right now, but I don't know how they simplified it to that..

OpenStudy (anonymous):

First partial fractions, then complete the square, and then get clever and look for the right inverse transforms on the table.

OpenStudy (anonymous):

Thanks! I figured it out. So the answer is \[3-2\sin2t+5\cos2t\] but i don't know how to get the 2sin2t..

OpenStudy (anonymous):

this is a partial fraction to solve the other part 5s-4/(s^2+4)+3s =5s-4/(s^2+3s+4) by factorisation the denominator gives (s-1)(s+4) therefore 5s-4/(s-1)(s+4) which in turns gives A/(s-1) +B/(s+4) and using cover up rule the for A 5s-4/s+4 as s tends to 1 we have 5(1)-4/1+4=1/5 and for B 5s-4/s-1 as s tends to -4 gives -20-4/-4-5=-24/5 the substituting in the above gives 1/5(s-1) -24/5(s+4)

OpenStudy (anonymous):

8s^2-4s+12/s(s^2+4) =A/s+Bs+C/s^2+4 using partial fraction and collecting co effiecient 8s^2-4s+12=A(s^2+4)+s(Bs+C) =As^2+4A+Bs^2+Cs collecting co efficient 8s^2=s^2(A+B) equation 1 -4s=Cs equation 2 12=4A equation 3 then solving for the values from equn 3 A=12/4=3 substituting values in equn 1 8=A+B 8=3+B then B=5 and C=-4 from equn 2 then putting back all the values we have 3/s+5-4/s^2+4